HUBBLE SPACE TELESCOPE DAILY REPORT #2565 PERIOD COVERED: 0000Z (UTC) 02/22/00 - 0000Z (UTC) 02/23/00 Daily Status Report as of 054/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was to be used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in HSTAR 7521 and yesterday's report, the acquisition/re-acquisitions for both iterations of this proposal failed and the take data flag remained down. 1.2 Completed WF/PC-2 8242 (Imaging of Dust Structures and Star Formation in the Disk-Halo Interface of Spiral Galaxies) The WF/PC-2 was to be used to obtain images of five edge-on spirals to study star formation and dusty interstellar clouds in the disk-halo interface of these galaxies. As described in HSTAR 7521 and yesterday's report, the acquisition/re-acquisitions for this proposal failed and the take data flag remained down. 1.3 Completed Four Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. As described in HSTAR 7521 and yesterday's report, the acquisition/re-acquisitions for the first two iterations of this proposal failed and the take data flag remained down. Otherwise, the observations were completed as planned, and no further nomalies were reported. 1.4 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was to be used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. As described in HSTAR 7521 and yesterday's report, the acquisition/re-acquisitions for this proposal failed and the take data flag remained down. 1.5 Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was to be used to monitor the darks. As described in HSTAR 7521 and yesterday's report, the acquisition/re-acquisitions for this proposal failed and the take data flag remained down. 1.6 Completed FGS 8527 (Near Term Stability) FGS-2R was used to determine the near term stability of the instrument. The goals are twofold: {1} to establish the baseline Long Term Stability measurement for the new FGS and monitor changes in the relative plate scale and distortion since the SMOV MiniOFAD; and {2} to monitor changes in the transfer function over months. See 2.3 and HSTAR 7522 for details of instances of failed astrometry during this proposal. 1.7 Completed Three Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.9 Completed STIS/MA1 8169 (HST-Galileo Io Campaign) The Space Telescope Imaging Spectrograph (MA1) was used to make observations of Jupiter's moon Io. Io plays a central role in the Jovian system, yet the mechanisms that couple its volcanoes, the tenuous atmosphere, the escaping corona and the plasma torus remain elusive. In 1999 there is the unique opportunity to combine HST observations with the Galileo close flybys of Io. The observations were completed as planned, and no problems were reported. 1.10 Completed STIS/CCD/MA1/MA2 8171 (HST Far-UV Imaging and Spectra of Jupiter's Aurora During the Galileo Extended Mission) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make far-ultraviolet spectrographic observations of the aurora on the planet Jupiter as part of the extended Galileo Mission. Observations were also made of the moon Europa. This is to study the nature of the satellite auroral footprints in Jupiter's atmosphere. The observations were completed as planned, and no problems were reported. 1.11 Completed STIS/CCD/MA1/MA2 8317 (Mass Accretion Rates for Pre-Main Sequence Intermediate Mass Stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to obtain the first direct, accurate measurements of mass accretion rates for intermediate mass {1.5M{sun} to 3.5M{sun}} T Tauri stars}, precursors of the Beta Pic- type stars. The observations completed nominally. 1.12 Completed WF/PC-2 8445 (WF/PC-2 Cycle 8 Earth Flats) The WF/PC-2 was used to make Cycle 8 Earth flat calibration observations. This is to obtain sequences of earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 8436 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to make archive observations parallel to the prime science observations as part of a POMS test proposal. This is a POMS test proposal designed to simulate scientific plans. The observations were completed as planned, and no anomalies were reported. 1.14 Completed Two Sets of STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ0827+5 and HCLQ1658+3 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 8459 (Decontaminations and Associated Observations Pt. 2/2) The WF/PC-2 was used to perform the monthly decons. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, and UV throughput checks. The observations were completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 13 Successful: 13 Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 28 Successful: 28 2.3 Operations Notes: Servicing Mission Observatory Verification (SMOV) is continuing. Per ROP DF-18A, an ESB dump was performed and 14 messages were cleared at 053/1310Z. At 053/095621Z, the FGS proposal detailed in 1.6 had a successful guide star acquisition. At 054/102046Z, following a ZOE and about halfway through the proposal, it was noticed that there had been indications of failed astrometry. These indications occurred five more times during the remainder of this approximate 50-minute proposal. A "G" string line outage recorder playback was completed at 054/0315Z. This was followed by a"G string re-cycle, when the string was configured for real-time backup. The string is up and functioning well, but the CCM box, which monitors string problems, is having problems. These will be addressed in the morning. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW